OpenStudy (anonymous):

Solve for b1. a=(b1+b2)h/2 b1 = 2a/b2h b1 = 2a/h - b2 b1 = 2A - b2h

OpenStudy (anonymous):

the second one

Parth (parthkohli):

Yes. second is the correct answer.

OpenStudy (zzr0ck3r):

first multiply both sides by 2 2a=(b1+b2)h then devide by h 2a/h=b1+b2 then subtract b2 from both sides 2a/h-b2=b1